Write an equation in slope-intercept form for the line that is parallel to the given line and that passes through the given point. 5x - 2y = 10; (3, -5)
Answer by robertb(4012)

The equation should be of the form 5x - 2y = c, where c is an unknown constant. To find it, use the coordinates of the given point: 5*3 - 2*-5 = 15 + 10 = 25. Therefore the equation is 5x - 2y = 25.